OpenStudy (anonymous):

state the domain and range of the relation: x^2 + y^2 = 4

OpenStudy (amistre64):

since its a circle, it tends to double back on itself; its centered at the origin and a radius of 2 so it can never go beyond 2 units from (0,0)

OpenStudy (anonymous):

How do you know it only has a radius of two?

OpenStudy (amistre64):

becasue the equation of a circle is: x^2 + y^2 = r^2 .... since r^2 = 4, im gonna assume the r = 2

OpenStudy (anonymous):

Oh ok. Thank you =]

OpenStudy (amistre64):

youre welcome :)